OpenStudy (camper4834):

This limit should be simple but I can't figure it out! \[\lim_{t\rightarrow0}~~\frac{t^2}{\sin^2(t)}\]

OpenStudy (anonymous):

What? \(\sin\) of what variable? Also the limit variable is \(n\)?

OpenStudy (camper4834):

sorry i guess i wrote it kinda silly

OpenStudy (camper4834):

\[\lim_{t\rightarrow0}~~\frac{t^2}{\sin^2(t)}\]

OpenStudy (camper4834):

so...

OpenStudy (anonymous):

I think the limit is 1. Use L'hopital rule and take the derivative of the top and the derivative of the bottom separately and then blog in 0 when you can't take the derivative anymore which them would equal to 2/2 which is 1.

OpenStudy (camper4834):

i did use lhopital and you get \[\frac{2t}{2\sin\cos}\]

OpenStudy (camper4834):

and 2sincos is just 2sin(2\(\theta\))

OpenStudy (camper4834):

and so im back to 0/0

OpenStudy (camper4834):

*facepalm* no you are right. i should have done lhopitals twice. But i got 2/4

OpenStudy (anonymous):

the denominator will be -sin(t)sin(t) + cos(t)cos(t) -sin(t)sin(t) + cos(t)cos(t) so when you blog in 0 it would equal to 0 + 1 - 0 + 1 so it would be 2.

OpenStudy (anonymous):

\[ \frac{t^2}{\sin^2(t)} = \left(\frac{\sin t}{t}\right)^{-2} \]Since we know the inner limit exists we can do this: \[ \lim_{t\to 0}\left(\frac{\sin t}{t}\right)^{-2}=\left(\lim_{t\to 0}\frac{\sin t}{t}\right)^{-2} = (1)^{-2} = 1 \]

OpenStudy (camper4834):

are you sure you can do that?

OpenStudy (camper4834):

ive never heard of this rule before

OpenStudy (anonymous):

Really, you haven't?

OpenStudy (anonymous):

What rules have you heard of?

OpenStudy (anonymous):

It's just the division rule twice.

OpenStudy (anonymous):

\[ \lim_{x\to a}f(x) = L \quad \text{and}\quad \lim_{x\to a}g(x) = K\implies \lim_{x\to a}\frac{f(x)}{g(x)} =\frac{L}{K} \]First let \(f(x) = 1\), which means \(\lim_{x\to a}\frac{1}{g(x)} =\frac{1}{K}\) Then let \(f(x) = 1/g(x)\), which means \(\lim_{x\to a}\frac{1}{{g(x)}^2} =\frac{\frac{1}{K}}{K} = \frac{1}{K^2}\)